# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ F.3 Additional theoretical results
$\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\mathrm{rank}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|.$
Proof of Lemma 5.
Notice that for any $\mathcal{G}\preceq\mathcal{G}_{1}$, we have that $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})$; otherwise, conditional dependencies of $\mathcal{G}$ are not a subset of those of $\mathcal{G}_{1}$. Similarly, we have that for any $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{2})$. This allows us to conclude that for $\mathcal{G}$ with $\mathcal{G}\preceq\mathcal{G}_{1}$ and $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})$. Since $\mathrm{rank}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result. ∎
Lemma 6.
@@
$\min_{\mathcal{G}_{1}\preceq\mathcal{G},\mathcal{G}_{2}\preceq\mathcal{G}}\mathrm{rank}(\mathcal{G})\geq|\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})|.$
Proof of Lemma 6.
Based on the same argument as proof of Lemma 5, we have that for $\mathcal{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}$ and $\mathcal{G}_{2}\preceq\mathcal{G}$, $\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})\subseteq\mathrm{sk}(\mathcal{G})$. Since $\mathrm{rank}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result. ∎
Appendix G Theoretical results for MPDAGs
G.1 Poset for MPDAGs is not graded: proof of Proposition 10
To prove that the poset $\mathcal{L}$ underlying MPDAGs is not graded, it suffices to find two maximal chains between a pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ of different lengths. Fig.˜9 presents a collection of graphs that specify a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. This chain has length $3$. Similarly, Fig.˜10 presents another collection of graphs that specify a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. This chain, however, has length $5$. As the length of these two maximal chains are not equal, we conclude that $\mathcal{L}$ is not graded.
Figure 9: A maximal chain between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, specified by the sequence $(\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4})$ where $\mathcal{G}_{i+1}$ covers $\mathcal{G}_{i+1}$ for $i=1,2,3$.
Figure 10: A maximal chain between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, specified by the sequence $(\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4},\mathcal{G}_{5},\mathcal{G}_{6})$ where $\mathcal{G}_{i+1}$ covers $\mathcal{G}_{i+1}$ for $i=1,2,3,4,5$.
G.2 MPDAG model complexity: proof of Proposition 12
We construct an explicit sequence of MPDAG from $\mathcal{G}$ to the empty MPDAG by first directing one undirected at a time in $\mathcal{G}$ to obtain a sequence $\mathcal{G}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}$ such that $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{i+1})=1,\mathcal{G}_{i}\succeq\mathcal{G}_{i+1}$, $\mathcal{G}_{k}$ is a DAG and where $k$ is the number of undirected edges in $\mathcal{G}$. This is possible by repeatedly applying the result from Lemma 7. We append to this sequence another sequence $\mathcal{G}_{k}=\tilde{\mathcal{G}}_{1},\dots,\tilde{\mathcal{G}}_{l}$ such that $\mathrm{pseudo}\text{-}\mathrm{rank}(\tilde{\mathcal{G}}_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(\tilde{\mathcal{G}}_{i+1})=1,\tilde{G}_{i}\succeq\tilde{G}_{i+1}$, all $\tilde{\mathcal{G}}_{i}$ are DAGs, $\tilde{\mathcal{G}}_{l}$ is the empty graph and $l$ is the number of directed edges in $\mathcal{G}$. We can obtain such a sequence simply by deleting the edges in $G_{k}$ in any random order. By combining the two sequence we obtain a $\mathrm{pseudo}\text{-}\mathrm{rank}(G)$-length sequence of graphs from $G$ to the empty graph such that adjacent graphs on the sequence neighbors in $\mathcal{L}$ while having pseudo-ranks that differ by $1$.
Lemma 7.
Consider an MPDAG $\mathcal{G}$. Then there exists an undirected edge in $\mathcal{G}$ such that directing the edge results in an MPDAG $\tilde{\mathcal{G}}$ which is equal to $\mathcal{G}$ except for the one additional directed edge.
Figure 11: Forbidden induced subgraphs of an MPDAG (left) and induced subgraphs that may exist in an MPDAG and are one directed edge away from a forbidden subgraph (right).
@@ G.2 MPDAG model complexity: proof of Proposition 12
Consider $\mathcal{G}$ and suppose there exists a pair $(a,b)$, such that aba-b is in $\mathcal{G}$, $\mathrm{Sib}(a)\subseteq\mathrm{Sib}(b)\cup\mathrm{Ch}(b)$ and $b$ is a minimal element in the partial topological ordering of $\mathcal{G}$ among all nodes with an adjacent undirected edge. Here, $Sib(v)$ is the set of nodes $u$ such that uvu-v called the sibling of $v$, while $Ch(v)$ is the set of nodes $v\rightarrow u$ called the set of children of $v$. We first show that directing the edge to $a\rightarrow b$ results in an MPDAG $\tilde{\mathcal{G}}$ before showing that there must always exist such a pair. To show that $\tilde{\mathcal{G}}$ is an MPDAG it suffices to show that it does not contain any of the forbidden subgraphs shown on the left hand size of Fig. 11. In order for that to possibly occur $\mathcal{G}$ must contain a pre-forbidden structure as shown on the right hand size of Fig. 11. We now prove that if such a pre-forbidden structure exists we will never direct an edge that will result in a forbidden structure.
Consider the first pre-forbidden subgraph. The node in the middle always has at least one sibling that the two adjacent nodes do not have. Therefore, the middle node can never be selected as $b$ and we will never arrive at the forbidden structure. Consider the second pre-forbidden subgraph. The left most node is not minimal in the topological order among all nodes with an adjacent undirected edge and can therefore not be selected as $b$. The right most-node has one more sibling (the left most node) than the middle node. Therefore, we can never arrive at the forbidden subgraph. There is no pre-forbidden graph for the third structure. For the fourth pre-forbidden graph the node at the bottom-left is not minimal and can therefore never be selected as $B$.
It now remains to show that such a pair $(a,b)$ exists in any MPDAG that has at least one undirected edge. Consider the CPDAG $\mathcal{C}$ of $\mathcal{G}$. The connected undirected components of $\mathcal{C}$ are chordal and therefore each have a perfect numbering, i.e., an ordering $v_{1},\dots,v_{p}$ such that each $v_{i}$ is simplical in the induced subgraph of $\mathcal{C}$ on $v_{i},\dots,v_{p}$. Here a node $v$ is simplical if it’s neighborhood $\mathrm{Sib}(v)$ forms a clique. Furthermore, each member of a connected undirected component of $\mathcal{C}$ is equivalent in terms of the graphs topological order (Andersson, Madigan and Perlman, 1997). Therefore, the simplical node $s$ in the undirected component that is minimal in terms of the order satisfies i) minimality among all nodes with a sibling and ii) for any sibling $s^{\prime}$ of $s$ it holds that $\mathrm{Sib}(s)\subseteq\mathrm{Sib}(s^{\prime})$. In any MPDAG in which $s$ remains minimal among all nodes with siblings, any node $s^{\prime}$ topologically maximal among the siblings of $s$ satisfies $\mathrm{Sib}(s)\subseteq\mathrm{Sib}(s^{\prime})\cup\mathrm{Ch}(s^{\prime})$.
It thus only remains to consider the case that $s$ is not topologically minimal in $\mathcal{G}$. This implies that there exists an edge out of $s$ in $\mathcal{G}$ into a node $c$ that is a sibling of $s$ in $\mathcal{C}$. In order to avoid any forbidden structure of the first kind any undirected edge adjacent to $c$ in $\mathcal{C}$ that is towards a non-sibling of $s$ in $\mathcal{C}$ must also be directed in $\mathcal{G}$. Therefore, any sibling of $c$ in $\mathcal{G}$ must have been a sibling of $s$ in $\mathcal{G}$. As the siblings of $s$ in $\mathcal{G}$ form a clique it follows that there must exist a sibling $s^{\prime}$ of $c$ such that $\mathrm{Sib}(c)\subseteq\mathrm{Sib}(s^{\prime})\cup\mathrm{Ch}(s^{\prime})$. If $c$ is minimal it therefore satisfies the conditions we make on $b$. If not we repeat the argument until we arrive at such a $b$ through breadth first search.
If there are no undirected edges adjacent to $s$ in $\mathcal{G}$ we move to the next simplical node in the perfect ordering of the undirected subgraph of $\mathcal{C}$ and repeat the argument above.
G.3 Pseudo-distance upper-bounding SHD: proof of Proposition 13
@@
The proof is based on the following observation: for any pair of graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{2})-\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{1})=1$, the graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are identical except for a change in a single edge. This change can be one of the following: a missing edge becomes a directed edge, or a directed edge becomes an undirected edge.

### v2

@@ F.2 Additional results
$\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\operatorname{\mathrm{rank}}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|.$
Proof of Lemma S6.
Notice that for any $\mathcal{G}\preceq\mathcal{G}_{1}$, we have that $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})$; otherwise, conditional dependencies of $\mathcal{G}$ are not a subset of those of $\mathcal{G}_{1}$. Similarly, we have that for any $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{2})$. This allows us to conclude that for $\mathcal{G}$ with $\mathcal{G}\preceq\mathcal{G}_{1}$ and $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})$. Since $\operatorname{\mathrm{rank}}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result. ∎
Lemma S7.
@@
$\min_{\mathcal{G}_{1}\preceq\mathcal{G},\mathcal{G}_{2}\preceq\mathcal{G}}\operatorname{\mathrm{rank}}(\mathcal{G})\geq|\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})|.$
Proof of Lemma S7.
Based on the same argument as proof of Lemma S6, we have that for $\mathcal{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}$ and $\mathcal{G}_{2}\preceq\mathcal{G}$, $\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})\subseteq\mathrm{sk}(\mathcal{G})$. Since $\operatorname{\mathrm{rank}}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result. ∎
Appendix G Further results for MPDAGs
G.1 MPDAG model complexity: proof of Proposition 10
We will use the following notation and terminology in this subsection. The siblings of a vertex $v$ in graph $\mathcal{G}$, denoted by $\operatorname{Sib}_{\mathcal{G}}(v)$, are those vertices that are adjacent to $v$ via an undirected edge. We use $\operatorname{Pa}_{\mathcal{G}}(v)$ and $\operatorname{Ch}_{\mathcal{G}}(v)$ to denote the parents and the children of $v$ in $\mathcal{G}$ respectively. The neighbors of $v$ in $\mathcal{G}$, denoted by $\operatorname{N}_{\mathcal{G}}(v)$, is the set of vertices that are adjacent to $v$. By definition, it holds that $\operatorname{N}_{\mathcal{G}}(v)=\operatorname{Sib}_{\mathcal{G}}(v)\cup\operatorname{Pa}_{\mathcal{G}}(v)\cup\operatorname{Ch}_{\mathcal{G}}(v)$. The graph in the subscripts will be suppressed when it is clear from the context. A vertex set is called a clique if every pair of vertices in the set are adjacent. A vertex $v$ in $\mathcal{G}$ is called simplicial if $\operatorname{N}_{\mathcal{G}}(v)$ is a clique. The topological order is a partial order on the set of vertices: for two vertices $u$ and $v$, we say $u\preceq v$ if $u\rightarrow\dots\rightarrow v$. Given a vertex set $\mathrm{S}$ that contains a vertex $u$, we say $u$ is topologically minimal in $\mathrm{S}$ if there exists no vertex $v$ in $\mathrm{S}$ such that $v\prec u$.
We first prove a supporting lemma.
Lemma S8.
Let $\mathcal{H}$ be an MPDAG where every pair of vertices are connected by an undirected path. Suppose $u$ is a simplicial vertex and $u$ has a child $v$ with usvu-s-v in $\mathcal{H}$. Then $v$ is also simplicial.
[…]
Figure S11: Graphs that appear in the proof of Lemma˜S8.
@@ G.1 MPDAG model complexity: proof of Proposition 10
Because $u$ is simplicial, it suffices to show that $w\in\{u\}\cup\operatorname{N}(u)$ holds for every $w\in\operatorname{N}(v)$. For each $w\in\operatorname{N}(v)$, there are three cases as illustrated in Fig.˜S11. For each case, the argument is based on the four induced subgraphs listed in Fig.˜2(b) that are forbidden in any MPDAG.
  1. (I)
    $w\in\operatorname{Sib}(v)$. The result holds trivially if $w=s$. For a contradiction, suppose $w\neq s$ and $w\notin\operatorname{N}(u)$. Now observe that $u\rightarrow v-w$, as an induced subgraph, is forbidden.
  2. (II)
    $w\in\operatorname{Pa}(v)$. Suppose $u,w$ are not adjacent. To avoid forbidden structure $w\rightarrow v-s$, we have an edge between $s$ and $w$. Further, the edge cannot be undirected because otherwise the graph induced by $\{u,v,w,s\}$ is a forbidden structure. Therefore, the edge must be directed, in particular, as $w\rightarrow s$ due to the undirected edge svs-v. However, this induces another forbidden structure $w\rightarrow s-u$.
  3. (III)
    $w\in\operatorname{Ch}(v)$. Because every pair of vertices is connected by an undirected path in $\mathcal{H}$, let $p$ be a shortest undirected path that connects $w$ and $\{u,s,v\}$. We prove $w\in\operatorname{N}(u)$, in particular, $u\rightarrow w$, by induction on the length of $p$.
    1. (i)
      Base case: $p$ is of length 1. Observe that in this case we must have an undirected edge between $s$ and $w$. Without an edge between $u$ and $w$, the graph induced by $\{s,u,v,w\}$ is a forbidden structure. Hence, we have an edge between $u$ and $v$, in particular, $u\rightarrow w$ by acyclicity.
    2. (ii)
      Suppose the result holds when $p$ is of length $l$ or shorter for $l\geq 1$. Now we show that the result still holds when $p$ is of length $l+1$. We write the path as $p=\langle w\equiv w_{0},w_{1},\dots,w_{l+1}\rangle$ for $w_{l+1}\in\{s,u,v\}$. To avoid the forbidden structure $v\rightarrow w-w_{1}$, we know $v$ and $w_{1}$ are adjacent. There are two further cases.
      1. (a)
        $v\rightarrow w_{1}$. Observe that $w_{1}$ plays the role of $w$ and $w_{1}$ is connected to $\{s,u,v\}$ by an undirected path of length $l$. By our induction hypothesis, we have $u\rightarrow w_{1}$. To avoid the forbidden structure $u\rightarrow w_{1}-w$, there must be an edge between $u$ and $w$. In particular, we have $u\rightarrow w$.
      2. (b)
        $v-w_{1}$. By case (I), we know $u$ and $w_{1}$ are adjacent. To prove by contradiction, suppose there is no edge betweeb $u$ and $w$. Observe that the edge between $u$ and $w_{1}$ cannot be undirected since otherwise $\{u,v,w,w_{1}\}$ induces a forbidden structure. Further, due to $w_{1}-v$, we cannot have $w_{1}\rightarrow u$. Hence, we must have $u\rightarrow w_{1}$. However, we see a contradiction because this leads to a forbidden structure $u\rightarrow w_{1}-w$. Hence, there must be an edge between $u$ and $w$, in particular, $u\rightarrow w$.
Figure S12: Induced subgraphs that are forbidden in an MPDAG (left) and their corresponding pre-structures (right). Each pre-structure can exist as an induced subgraph in an MPDAG, but upon directing aba-b to $a\rightarrow b$, it becomes a forbidden structure.
We are ready to prove Lemma˜3.
Proof of Lemma˜3.
The proof consists of two parts. First, we show that by choosing an undirected edge aba-b that meets certain criteria (assuming such aba-b exists in $\mathcal{G}$) and directing it to $a\rightarrow b$ and leaving everything else intact, we obtain an MPDAG $\tilde{\mathcal{G}}$. Second, we argue that an edge aba-b that meets our selection criteria always exists in $\mathcal{G}$.
We show the first part. We select an edge aba-b in $\mathcal{G}$ that meets the following criteria:
  1. (i)
    $\operatorname{Sib}_{\mathcal{G}}(b)\subseteq\{a\}\cup\operatorname{Sib}_{\mathcal{G}}(a)\cup\operatorname{Ch}_{\mathcal{G}}(a)$, and
  2. (ii)
    there is no vertex $c$ such that $b\rightarrow c$ and bscb-s-c exist in $\mathcal{G}$.
Let $\tilde{\mathcal{G}}$ be a copy of $\mathcal{G}$ except that aba-b is directed to $a\rightarrow b$. We argue that $\tilde{\mathcal{G}}$ is an MPDAG by showing that it satisfies the graphical constraints listed in Section˜B.2. The chordality constraint is satisfied because $\mathcal{G}$ is already an MPDAG and we have only directed an undirected edge. Thus, it remains to show that the four forbidden structures listed in Fig.˜2(b) are not introduced. By inspecting these forbidden structures, we list in Fig.˜S12 the only possible ways that a forbidden structure can be introduced upon directing a single edge: forbidden structure III cannot be introduced; for structures I, II and IV to be introduced, a corresponding pre-forbidden structure must exist in $\mathcal{G}$ and aba-b must be chosen as specified in the figure. To finish proving the first part, we argue that our selection criteria prevent aba-b from being chosen in such ways:
  • Structure I.
    This violates (i);
  • Structure II.
    The first case violates (i) and the second case violates (ii);
  • Structure IV.
    This violates (ii).
We now prove the second part, namely the existence of aba-b in $\mathcal{G}$ that meets our selection criteria. By our assumption, $\mathcal{G}$ has one or more undirected edge. Accordingly, let $\mathrm{C}$ be a (maximal) connected component spanned by undirected edges that has two or more vertices and let $\mathcal{H}:=\mathcal{G}_{\mathrm{C}}$.
We first choose $b$ in $\mathcal{H}$ that satisfies criterion (ii). By Section˜B.2, $\mathcal{H}$ is an MPDAG and the skeleton of $\mathcal{H}$ is chordal. By a classical result on chordal graphs (Lauritzen, 1996, Lemma 2.9), there exists a simplicial vertex $b^{\prime}$ in $\mathcal{H}$ with $\operatorname{Sib}_{\mathcal{H}}(b)\neq\emptyset$. If $b^{\prime}$ satisfies (ii), let $b:=b^{\prime}$. Otherwise, whenever $b^{\prime}\rightarrow c$ and $b^{\prime}-s-c$ holds in $\mathcal{G}$ (and hence in $\mathcal{H}$ by definition), replace $b^{\prime}$ with $c$. We repeat this process until (ii) is satisfied and let $b$ be the final $b^{\prime}$. Because $\mathcal{H}$ has no directed cycle, the process must terminate in a finite number of steps. Further, by Lemma˜S8, the process maintains the property that $b^{\prime}$ is simplicial. Meanwhile, the fact that $\operatorname{Sib}_{\mathcal{H}}(b^{\prime})\neq\emptyset$ is also maintained. Hence, the resulting $b$ is simplicial in $\mathcal{H}$ with $\operatorname{Sib}_{\mathcal{H}}(b)\neq\emptyset$.
Finally, we choose $a\in\operatorname{Sib}_{\mathcal{H}}(b)$ such that $a$ is topologically minimal in $\operatorname{Sib}_{\mathcal{H}}(b)$. It only remains to show that criterion (i) holds. Since $b$ is simplicial in $\mathcal{H}$, for each $u\in\operatorname{Sib}_{\mathcal{H}}(b)$, we have $u=a$ or $u\in\mathcal{N}_{\mathcal{H}}(a)$. Further, by the way that $a$ is chosen, we know $u\notin\operatorname{Pa}_{\mathcal{H}}(a)$. Hence, we have $u\in\{a\}\cup\operatorname{Sib}_{\mathcal{H}}(a)\cup\operatorname{Ch}_{\mathcal{H}}(a)$, i.e.,
$\operatorname{Sib}_{\mathcal{H}}(b)\subseteq\{a\}\cup\operatorname{Sib}_{\mathcal{H}}(a)\cup\operatorname{Ch}_{\mathcal{H}}(a).$
By the fact that $\operatorname{Sib}_{\mathcal{G}}(b)=\operatorname{Sib}_{\mathcal{H}}(b)$, $\operatorname{Sib}_{\mathcal{G}}(a)=\operatorname{Sib}_{\mathcal{H}}(a)$ and $\operatorname{Ch}_{\mathcal{H}}(a)\subseteq\operatorname{Ch}_{\mathcal{G}}(a)$, we obtain
$\operatorname{Sib}_{\mathcal{G}}(b)\subseteq\{a\}\cup\operatorname{Sib}_{\mathcal{G}}(a)\cup\operatorname{Ch}_{\mathcal{G}}(a),$
i.e., criterion (i). ∎
Proof of Proposition˜10.
Suppose $\mathcal{G}$ has $l$ undirected edges and $k$ directed edges, so we have $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G})=2l+k$. The maximal chain is given by
$\mathcal{G}=:\mathcal{G}_{2l+k}\gtrdot\mathcal{G}_{2(l-1)+k+1}\gtrdot\dots\gtrdot\mathcal{G}_{l+k}\gtrdot\mathcal{G}_{l+k-1}\gtrdot\dots\gtrdot\mathcal{G}_{0}:=\mathcal{G}_{\hat{0}}$,
which consists of two segments: (i) starting from $\mathcal{G}$, we repeatedly apply Lemma˜3 to convert one undirected edge to a directed edge until we reach $\mathcal{G}_{l+k}$ that only contains directed edges; (ii) then, we remove directed edges iteratively in any order till we reach the empty graph. ∎
G.2 Pseudo-distance upper-bounding SHD: proof of Proposition 11
@@
The proof is based on the following observation: for any pair of neighboring graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})-\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})=1$, the graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are identical except for a single edge: either a missing edge versus a directed edge, or a directed edge versus an undirected edge.